# manvr3d: A Platform for Human-in-the-loop Cell Tracking in Virtual Reality

### v1

@@ 5 Performance
We benchmarked sciview’s capability to render geometry instances (see Table 1) with two differently-sized datasets and found that for up to several ten thousand cells, there is no negative effect on performance [^4] The frame rates refer to a render resolution of 1920x1080 pixels, on an XMG Fusion 15 laptop with NVIDIA RTX 4070 GPU and Intel Core i9-194900HX processor, running Windows 10 build 19044.5371.. Turning off the overlayed volume rendering yields slightly faster frame rates.
@@ Acknowledgements.
The authors thank Vladimír Ulman and Ruoshan Lan for their work on a prototype of this project. Thanks to Elias Barriga, Valentin Ruffine, Ana Patricia Ramos and Jaime Hidalgo for their feedback on the software. This work was partially funded by the Center for Advanced Systems Understanding (CASUS), financed by Germany’s Federal Ministry of Education and Research (BMBF) and by the Saxon Ministry for Science, Culture and Tourism (SMWK) with tax funds on the basis of the budget approved by the Saxon State Parliament. This work was also funded by the French National Research Agency (France BioImaging, ANR-24-INBS-0005 FBI BIOGEN).

### v2

@@ 5 Performance
We benchmarked sciview’s capability to render geometry instances (see Table 1) with two differently-sized datasets and found that for up to several ten thousand cells, there is no negative effect on performance [^4] The frame rates refer to a render resolution of 1920x1080 pixels, on an XMG Fusion 15 laptop with NVIDIA RTX 4070 GPU and Intel Core i9-14900HX processor, running Windows 10 build 19044.5371, JDK version 21.0.5, NVIDIA driver version 566.36.. Turning off the overlayed volume rendering yields slightly faster frame rates.
@@ Acknowledgements.
The authors thank Ko Sugawara for his ongoing support with the ELEPHANT integration. Thanks to Vladimír Ulman and Ruoshan Lan for their work on a prototype of this project. The authors also want to thank Elias Barriga, Valentin Ruffine, Ana Patricia Ramos and Jaime Hidalgo for their feedback on the software. This work was partially funded by the Center for Advanced Systems Understanding (CASUS), financed by Germany’s Federal Ministry of Education and Research (BMBF) and by the Saxon Ministry for Science, Culture and Tourism (SMWK) with tax funds on the basis of the budget approved by the Saxon State Parliament. This work was also funded by the French National Research Agency (France BioImaging, ANR-24-INBS-0005 FBI BIOGEN).